Supplemental Operations Report – March 23rd

Supplemental Operations Report 23 Mar 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Outpost water supply, power system

Action taken for non-nominal systems: Empty, clean and refill of water tank will take place tomorrow.

Generator: Still being manually run at night.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night. Unexplainable drop in SOC today.

Solar— Nominal

Notes on power system: At 7:15 tonight the SOC was 0% and the data show a continuous and steady power drain all day. It makes no sense because it was a sunny day and we did not plug in the rovers after use. I will troubleshoot it tomorrow but in the meantime, we might lose power tomorrow because we can’t charge the batteries unless we turn off the generator. This was a shock to me because I don’t know what could have caused this.

Propane Reading, station tank – not noted %
Propane Reading, director tank— not noted %
Propane Reading, intern tank— not noted %
Propane Reading, generator— not noted %

Ethanol Free Gasoline – 2 gallons.

Water (Outpost tank) – 450 gallons (currently being treated with bleach

Sojourner rover used: yes
Hours: not noted
Beginning charge: 100
Ending charge: 84
Currently charging: no

Notes on rovers: The crew is using all rovers, including Soji

ATV’s Used: (Honda, 350.1, 350.2, 300): Honda and the 350 that runs

Reason for use: crew use

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: not known hours

Notes on ATVs: nothing to report

HabCar used and why, where? Yes, crew is using it

CrewCar used and why, where? No, it is parked in Grand Junction

Luna used and why, where? Yes, to town

Campus-wide inspection, if action taken, what and why:

General notes and comments: Things are very chaotic and I’m not able to keep track of what is happening.

Summary of general operations: I replaced the section of broken conduit for the power to the Outpost. Still working on ways to protect it from being run over again.

Summary of internet: Nominal

Summary of suits and radios: Nothing to report

Summary of Hab operations: At some point on Monday afternoon, the water pump did not stop running and the crew could not find out the reason. In addition, they smelled propane yesterday and investigated but no indication of a leak on both meters and the alarm.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Used to make replacement parts for the broken conduit.

Summary of any observatory issues: On Tuesday the cords on the mount were reconfigured to allow freedom of movement
Summary of health and safety issues: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site